OpenStudy (anonymous):

Which is an example of the commutative property? A. –13 + 27 = 27 + (–13) B. –21 + (11 + 13) = (–21 + 11) + 13 C. 4(–3 + 5) = 4(–3) + 4 ∙ 5 D. –(–8 + 3) = 8 + (–3)

OpenStudy (anonymous):

MEDAL!

OpenStudy (anonymous):

Thank You

OpenStudy (airyana1114):

Wait, is this for addition or multiplication?

OpenStudy (anonymous):

Addition and Sub

Mehek (mehek14):

\(\bf{a+b=b+a}\) that's the rule

OpenStudy (airyana1114):

What do you think the best answer is seeing that rule?

OpenStudy (anonymous):

Not really understanding it yet.

Mehek (mehek14):

look for the answer choices where the numbers are in different order ex: \(\bf{3+5=5+3}\)

OpenStudy (airyana1114):

Basically whatever is on the left of the equal sign will be the same as on the right side just flipped.

OpenStudy (anonymous):

Is it D?

OpenStudy (airyana1114):

No

Mehek (mehek14):

no that is distributive

Mehek (mehek14):

same for C D and C are both distributive

OpenStudy (anonymous):

Oh, A?

Mehek (mehek14):

yes ^_^

OpenStudy (airyana1114):

Good Job :)

Mehek (mehek14):

the numbers were flipped but when you add, you get the same answer

OpenStudy (anonymous):

YaY :)